# Generalization in Reinforcement Learning for Radio Access Networks

### v2

@@ 8.4 Reward Analysis
Figures 8 and 9, together with the reward definition in Equation 2, highlight the dual role of the robustness weight $\alpha$ in the design of RL-based link adaptation.
Specifically, we trained three generalized policies using the same model architecture, hyperparameters, and training procedure as in the previous section, but with different settings of $\alpha$: 0, 0.5, and 2. Figure 8 shows how $\alpha$ affects the performance achieved by each policy when tested in the five benchmarks, presenting statistics of gains/losses in throughput (see Figure 8) and spectral efficiency (see Figure 8) relative to the OLLA baseline, as well as the average BLER (see Figure 8) and average MCS index selected (see Figure 8), with 95% confidence intervals. Moreover, Figures 9 and 9 further examine how $\alpha$ influences the learned policy by showing the cumulative distribution function (CDF) of the selected MCS index (i.e., the action distribution) and the corresponding BLER, respectively, for benchmark B2.
@@
Figure 9: Effect of robustness weight $\alpha$ in the reward function on the learned RL policy, illustrated through action distributions (MCS index) and BLER for three polices evaluated in benchmark scenario B2 from Figure 8. Increasing $\alpha$ from 0 to 2 encourages more conservative MCS selection, resulting in smaller data payloads better matched to link capacity, thereby reducing BLER distribution.
Setting $\boldsymbol{\alpha=0}$ deactivates the penalty term for failed transmissions (i.e., NACK receptions). This results in a policy that encourages the selection of more aggressive (i.e., higher) MCS indices, as shown in Figure 9, thereby increasing the raw symbol rate. Consequently, the policy learned with $\alpha=0$ tends to maximize spectral efficiency, as illustrated in Figure 8 across all benchmarks, but at the expense of lower throughput, as depicted in Figure 8. Specifically, the policy exploits retransmissions to deliver more information bits than the channel can support in a single transmission, resulting in higher BLER (see Figure 9), which ultimately degrades throughput.
Setting $\boldsymbol{\alpha}$ too high (e.g., $\boldsymbol{\alpha=2}$) has the opposite effect, as it heavily penalizes failed transmissions. The resulting policy promotes very conservative MCS selection, as shown in Figure 9 (in green). Selecting overly low MCS indices leads to transmissions with lower spectral efficiency that carry fewer information bits—but with high reliability. As a result, the policy improves transmission reliability, as evident from the lower BLER distribution in Figure 9, at the cost of reduced throughput and spectral efficiency across all benchmarks, as shown in Figures 8 and 8, respectively.
@@
  • Spectral efficiency decreases monotonically with increasing $\alpha$, particularly in high-capacity scenarios such as B4, where it drops from approximately $+40\%$ at $\alpha=0$ to about $+15\%$ at $\alpha=2$.
@@
Setting $\boldsymbol{\alpha}$ to optimize throughput requires a moderate penalty that balances spectral efficiency and reliability. In our evaluations, this balance occurs near $\alpha=0.5$. At this value, the policy reduces retransmissions without overly suppressing MCS aggressiveness, thereby improving throughput. This behavior is observed in most multi-cell scenarios from Table 2—e.g., B2(a/b), B3, B4, and B5—as well as in the interference-free scenario B1, where performance losses at $\alpha=0$ become modest gains at $\alpha=0.5$. Hence, the throughput curve exhibits a concave profile, peaking near $\alpha\approx 0.5$.
@@
Figure 11: Comparison between MLP-only and GNN-based models (GCN + MLP and GAT + MLP, as per Figure 10) for two versions of benchmark B2(b). Within each plot, the left-most group presents results for 1-site 1-cell version, while the right-most group refers to a denser 3-site 9-cell deployment. Images from left to right show average downlink cell throughput and spectral efficiency gain with respect to OLLA, as well as BLER, respectively.
@@ 8.5 Impact of Graph-Based Approach
We next analyze how using graph representations to encode static and semi-static network information—capturing both site/cell attributes and inter-site/cell relationships within a RAN topology—can improve the robustness and transferability of RL models across diverse deployment scenarios. Our findings, presented in Figure 11, indicate that attention-based graphs provide benefits even in relatively small network deployments (e.g., 9 cells), although the gains diminish in very small deployments (e.g., 3 cells), where there is insufficient structural diversity to exploit.
@@
The results in Figure 11 are obtained by encoding the static and semi-static network information outlined in Section 7.4.1, such as cell type, number of antennas, downlink transmit power, and average cell utilization, using a GNN. The overall model architecture, illustrated in Figure 10, consists of a cascade of a GNN followed by an MLP for Q-value estimation. The MLP receives as input both the encoding produced by the GNN and the raw features representing the fast dynamics of link adaptation.
@@
As introduced in Section 4, the input to the GNN consists of attribute vectors $\mathbf{h}_{i}$ for nodes $i$ in a graph representing the local network deployment surrounding a UE, where each node corresponds to either the serving cell or an interfering cell. The full set of node features is denoted by the matrix $\mathbf{H}^{(0)}$, where each row corresponds to the attribute vector of a node, and $\mathbf{H}^{(l)}$ represents the encoding produced at layer $l$ of the GNN.
@@
We consider two GNN-based architectures: GCN + MLP and GAT + MLP. In both cases, the GNN and subsequent MLP are trained end-to-end using backpropagation, minimizing Q-value prediction error following the procedure described in Section 8.1. While we tested all five benchmarks from Table 2, Figure 11 presents results only for benchmark B2(b), as similar conclusions were observed across benchmarks. For this analysis, we also include an extended version of B2 with three sites and nine cells.
@@
We compare these graph-based models with the MLP-only architecture introduced in Section 8.3, which consists of a 6-layer MLP with 256 neurons per layer and no GNN-based encoding. All models are trained with a robustness weight of $\alpha=0.5$ in Equation 2, and we use GNNs with a single layer due to the fully connected graph topology and to reduce computational complexity.
@@
Figure 11 presents results for all approaches in terms of: (a) throughput gain relative to the OLLA baseline, (b) spectral efficiency gain relative to the OLLA baseline, and (c) average experienced BLER. We observe that, while GNN-based and MLP-only models exhibit similar performance in small deployments (e.g., B2(b)), the GAT model achieves approximately 30% higher throughput than both the GNN-based and MLP-only models in the extended 9-cell deployment. This suggests that incorporating attention mechanisms, such as those in the GAT architecture, becomes increasingly beneficial in larger network surrounding, where the topology and interrelations between radio sites/cells become more complex and diverse.
@@ 9 CONCLUSIONS AND FUTURE EXTENSIONS
This work explored a critical challenge in the application of deep RL in radio access networks: achieving robust generalization across diverse and dynamic wireless environments. While RL holds promise for automating complex RAN control functions, such as those involved in RRM and network orchestration, its adoption in real-world RANs can be hindered when models trained in specific conditions fail to transfer across the heterogeneous deployment conditions, user behavior, channel dynamics, and so forth.
To address this challenge, we introduced a comprehensive framework to improve generalization in RL-based RAN applications. We began by formally defining RL generalization in the context of RANs, identifying key obstacles such as partial observability and the diversity and non-stationarity of radio environments. We then proposed three core enablers for the development of generalizable RL policies for RAN applications: improved state abstraction tailored to RAN-specific characteristics; diversification of training environments to expose models to a wide range of deployment scenarios and radio environment conditions; and a scalable architecture for distributed learning in RAN to support large-scale data collection across many simulated conditions or real network nodes. Our proposed AI architecture for RAN not only supports sim2sim generalization but also aligns with the structural and operational requirements of real-world RANs, making it suitable for deployment in live networks.
We validated the feasibility and advantages of our approach by designing a novel RL algorithm for link adaptation. We demonstrated, using a high-fidelity system-level simulator compliant with 5G NR systems, that RL algorithms trained with our generalization enablers can outperform fine-tuned models as well as state-of-the-art heuristics when evaluated across diverse and unseen environments. These results highlight the potential of deploying zero-shot generalization in realistic network settings, avoiding frequent retraining.
Overall, this work provides a concrete step toward scalable, generalizable AI solutions for wireless networks. By enabling policy robustness and simplifying lifecycle management, it paves the way for the practical integration of RL solutions into next-generation AI-native RANs.

### v3

@@ 9.4 Reward Analysis
Unlike the OLLA baseline, which operates under a 10% BLER target that can be reliably achieved with full-buffer traffic (e.g., benchmark B2), the proposed RL-based design has no explicit access to or control over BLER during training or inference. Figures˜8 and 9, together with the reward definition in Equation˜2, illustrate the dual role of the robustness weight $\alpha$ in the RL-based LA design, enabling the learning of policies that realize different tradeoffs among throughput, spectral efficiency, and BLER.
Specifically, we trained three generalized policies using the same model architecture, hyperparameters, and training procedure as in the previous section, but with different settings of $\alpha$: 0, 0.5, and 2. Figure˜8 shows how $\alpha$ affects the performance achieved by each policy when tested in the five benchmarks, presenting statistics of gains/losses in throughput (see Figure˜8) and spectral efficiency (see Figure˜8) relative to the OLLA baseline, as well as the average BLER (see Figure˜8) and average MCS index selected (see Figure˜8), with 95% confidence intervals. Moreover, Figures˜9 and 9 further examine how $\alpha$ influences the learned policy by showing the cumulative distribution function (CDF) of the selected MCS index (i.e., the action distribution) and the corresponding BLER, respectively, for benchmark B2.
@@
Figure 9: Effect of robustness weight $\alpha$ in the reward function on the learned RL policy, illustrated through action distributions (MCS index) and BLER for three polices evaluated in benchmark scenario B2 from Figure˜8. Increasing $\alpha$ from 0 to 2 encourages more conservative MCS selection, resulting in smaller data payloads better matched to link capacity, thereby reducing BLER distribution.
Setting $\boldsymbol{\alpha=0}$ deactivates the penalty term for failed transmissions (i.e., NACK receptions). This results in a policy that encourages the selection of more aggressive (i.e., higher) MCS indices, as shown in Figure˜9, thereby increasing the raw symbol rate. Consequently, the policy learned with $\alpha=0$ tends to maximize spectral efficiency, as illustrated in Figure˜8 across all benchmarks, but at the expense of lower throughput, as depicted in Figure˜8. Specifically, the policy exploits retransmissions to deliver more information bits than the channel can support in a single transmission, resulting in high BLER (see Figure˜9), which ultimately degrades throughput.
Setting $\boldsymbol{\alpha}$ too high (e.g., $\boldsymbol{\alpha=2}$) has the opposite effect, as it heavily penalizes failed transmissions. The resulting policy promotes very conservative MCS selection, as shown in Figure˜9 (in green). Selecting overly low MCS indices leads to transmissions targeting a spectral efficiency lower than the effective channel capacity, thus carrying fewer information bits—but with high reliability. As a result, the policy improves transmission reliability, as evident from the lower BLER distribution in Figure˜9, at the cost of reduced throughput and spectral efficiency across all benchmarks, as shown in Figures˜8 and 8, respectively.
@@
  • Spectral efficiency decreases monotonically with increasing $\alpha$, particularly in high-capacity scenarios such as B4, where it drops from approximately $40\%$ at $\alpha=0$ to about $15\%$ at $\alpha=2$.
@@
Setting $\boldsymbol{\alpha}$ to optimize throughput requires a moderate penalty that balances spectral efficiency and reliability. In our evaluations, this balance occurs near $\alpha=0.5$. At this value, the policy reduces retransmissions without overly suppressing MCS aggressiveness, thereby improving throughput. This behavior is observed in most multi-cell scenarios from Table˜2—e.g., B2(a/b), B3, B4, and B5—as well as in benchmark B1, where performance losses at $\alpha=0$ become modest gains at $\alpha=0.5$. Hence, the throughput curve exhibits a concave profile, peaking near $\alpha\approx 0.5$.
@@
Figure 11: Comparison between MLP-only and GNN-based models (GCN + MLP and GAT + MLP, as per Figure˜10) for two versions of benchmark B2(b). Within each plot, the left-most group presents results for 1-site 1-cell version, while the right-most group refers to a denser 3-site 9-cell deployment. Images from left to right show average downlink cell throughput and spectral efficiency gain with respect to OLLA, as well as BLER, respectively.
@@ 9.5 Impact of Graph-Based Approach
We next analyze how using graph representations to encode static and semi-static network information—capturing both site/cell attributes and inter-site/cell relationships within a RAN topology—can improve the robustness and transferability of RL models across diverse deployment scenarios. Our findings, presented in Figure˜11, indicate that attention-based graphs provide benefits even in relatively small network deployments (e.g., 9 cells), although the gains diminish in very small deployments (e.g., 3 cells), where there is insufficient structural diversity to exploit.
@@
The results in Figure˜11 are obtained by encoding the static and semi-static network information outlined in Section˜8.4.1, such as cell type, number of antennas, downlink transmit power, and average cell utilization, using a GNN. The overall model architecture, illustrated in Figure˜10, consists of a cascade of a GNN followed by an MLP for Q-value estimation. The MLP receives as input both the encoding produced by the GNN and the raw features representing the fast dynamics of link adaptation.
@@
As introduced in Section˜4, the input to the GNN consists of attribute vectors $\mathbf{h}_{i}$ for nodes $i$ in a graph representing the local network deployment surrounding a UE, where each node corresponds to either the serving cell or an interfering cell. The full set of node features is denoted by the matrix $\mathbf{H}^{(0)}$, where each row corresponds to the attribute vector of a node, and $\mathbf{H}^{(l)}$ represents the encoding produced at layer $l$ of the GNN.
@@
We consider two GNN-based architectures: GCN + MLP and GAT + MLP. In both cases, the GNN and subsequent MLP are trained end-to-end using backpropagation, minimizing Q-value prediction error following the procedure described in Section˜9.1. While we tested all five benchmarks from Table˜2, Figure˜11 presents results only for benchmark B2(b), as similar conclusions were observed across benchmarks. For this analysis, we also include an extended version of B2 with three sites and nine cells.
@@
We compare these graph-based models with the MLP-only architecture introduced in Section˜9.3, which consists of a 6-layer MLP with 256 neurons per layer and no GNN-based encoding. All models are trained with a robustness weight of $\alpha=0.5$ in Equation˜2, and we use GNNs with a single layer due to the fully connected graph topology and to reduce computational complexity.
@@
Figure˜11 presents results for all approaches in terms of: (a) throughput gain relative to the OLLA baseline, (b) spectral efficiency gain relative to the OLLA baseline, and (c) average experienced BLER. We observe that, while GNN-based and MLP-only models exhibit similar performance in small deployments (e.g., B2(b)), the GAT model achieves approximately 30% higher throughput than both the GNN-based and MLP-only models in the extended 9-cell deployment. This suggests that incorporating attention mechanisms, such as those in the GAT architecture, becomes increasingly beneficial in larger network surrounding, where the topology and interrelations between radio sites/cells become more complex and diverse.
@@ 10 Conclusions and Future Extensions
This work addressed a fundamental challenge in applying deep RL to RANs: achieving robust generalization across diverse and dynamic wireless environments. Although RL shows great promise for automating complex RAN control functions—such as those involved in RRM and network orchestration—its deployment in real-world RANs remains limited by the difficulty of transferring models trained under specific conditions to heterogeneous environments characterized by varying deployment scenarios, user behaviors, and channel dynamics.
To address this limitation, we introduced a comprehensive framework to improve generalization in RL-based RAN applications. We formally defined RL generalization in the context of RANs and identified key challenges, including partial observability and the diversity and non-stationarity of radio environments. We then proposed three core enablers for developing generalizable RL policies: (i) improved state abstraction tailored to RAN-specific characteristics; (ii) diversification of training environments to expose models to a wide range of deployment scenarios and radio environment conditions; and (iii) a scalable architecture for distributed learning, supporting large-scale data collection across simulated and real network nodes.
The proposed AI architecture supports not only sim2sim generalization but also alignment with the structural and operational requirements of real-world RANs, making it suitable for deployment in live networks. This capability enables models to be trained directly on field data collected across the network, thereby substantially mitigating performance discrepancies induced by the sim2real gap. In this setting, models initially trained in simulation can serve as a warm-up phase before fine-tuning on real-world data gathered from distributed network nodes.
We validated the feasibility and effectiveness of the proposed approach by developing a novel RL algorithm for link adaptation. Using a high-fidelity system-level simulator compliant with 5G NR standards, we demonstrated that RL algorithms trained with the proposed generalization enablers outperform both fine-tuned models and state-of-the-art heuristics when evaluated across diverse and previously unseen environments. These results highlight the potential of zero-shot generalization in realistic network settings, thereby reducing the need for frequent retraining.
Overall, this work presents a concrete step toward scalable and generalizable AI solutions for wireless networks. By enhancing policy robustness and simplifying lifecycle management, it paves the way for the practical integration of RL solutions into next-generation AI-native RANs.